OpenStudy (anonymous):

Please help. I tried doing it on my own its to confusing! give me a sec to post it. :)

OpenStudy (anonymous):

\[9\sqrt[4]{10+2}\sqrt[3]{10-6}\sqrt[4]{10-6}\sqrt[3]{10}\]

OpenStudy (anonymous):

it needs to be simplfied